OpenStudy (anonymous):

]Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = ; y intercept at (0, 1) slope = ; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

First, we pu the equation into the slope intercept-form: y = mx + b where m is the slope and b is the y-intercept y = (1/4)x - 1 so slope is 1/4 and y-intercept is -1 which is (0, -1)

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

The slope is (1/4, which is m) The y-intercept is (-1, which is b)

OpenStudy (anonymous):

Your choices are missing some characters. You have to go back to where you got the choices and find the one which matches the information I gave you.

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

i just copied it,

OpenStudy (anonymous):

Notice how your choices b and c have nothing after the slope, so they are incomplete as written.

OpenStudy (anonymous):

it has to be c becuase the choices with 4 cant b correct so the other one with y =-1 is c

OpenStudy (anonymous):

Then it is probably c because that's the only one with the right intercept and a non-misleading slope.

OpenStudy (anonymous):

yea because you probably used a picture as the fraction so you cant copy it

OpenStudy (anonymous):

So I would just go with c.

OpenStudy (anonymous):

nope, no picture.

OpenStudy (anonymous):

Sometimes the characters don't copy over. You'll have no problem if you go with c.

OpenStudy (anonymous):

yea its definately c

OpenStudy (anonymous):

But if you want to know how it's done, just look at my first post.

OpenStudy (anonymous):

]Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

That first post of mine shows you how to match up "m" with slope and "b" with the y-intercept. That slope-intercept form of the line is a very handy formula for your current and further math.

OpenStudy (anonymous):

y = mx+b y = -6x+b ... Plug in the given slope -3 = -6(-2)+b ... Plug in the given point

OpenStudy (anonymous):

b = -15

OpenStudy (anonymous):

The best way to solve this is to use a different form for the line, the point-slope formula, otherwise you are working backwards and harder. \[y - y _{1} = m(x - x _{1)}\]

OpenStudy (anonymous):

Here, you put in your point:\[(x _{1}, y _{1})\]which is: (-3, -1) So, y - (-1) = 4(x - (-3)) y + 1 = 4(x + 3)

OpenStudy (anonymous):

lol whoops i did the wrong problem lol my bad

OpenStudy (anonymous):

So, your choice is the second equation for this problem. This point-slope equation is another one good to memorize

OpenStudy (anonymous):

lol and thank you guys♥

OpenStudy (anonymous):

You are quite welcome.

OpenStudy (anonymous):

thx again, @ambermarie

OpenStudy (anonymous):

No, thank you :)

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

:)